# Hardwired-Neurons Language Processing Units as General-Purpose Cognitive Substrates

### v1

@@ 6.1. Hardware and System-Level Evaluation
We implemented the core components of HNLPU architecture, including HN Array, Control Unit, VEX, Interconnection Engine and on-chip Attention Buffer in RTL with Verilog, and verify the correctness of the RTL design through extensive test cases. We followed a standard ASIC design flow to obtain physical characteristics. The design was synthesized using Synopsys Design Compiler and placed-and-routed using Synopsys IC Compiler, on 5 nm technology. Power consumption was analyzed by PrimeTime PX using workload-derived switching activity (SAIF file) to accurately model both static and dynamic power. On-chip SRAMs were generated and analyzed using Memory Compiler on the same technology node.
@@ 6.2. Model
We selected the OpenAI gpt-oss 120 B model for system-level evaluation. It is a state-of-the-art open-source MoE large language model built on Llama-style architecture. We used 4-bit quantized version of the model and hardwiring the weights in HN Array. The HNLPU implementation of the model follows partitioning method, dataflow and mapping strategies detailed in Section 4 and Section 5.
@@ 6.3. Baseline Configurations
    Cerebras WSE-3: The throughput was empirically measured through publicly accessible Cerebras cloud service (cer, 2024a) running the gpt-oss 120 B model. As power measurement on cloud is not practical, we adopted the system power figures reported in (Kundu et al., 2025) instead.
@@
    HNLPU: The full system modeled as previous described.
@@ 7.1. Layout Characteristics
Table  1 presents the area and power breakdown of a single chip in the HNLPU. The HN Array and the Attention Buffer are the dominant components in terms of both area and power. The chip occupies a total area of 827.08 mm2 and has a power consumption of 308.39 W.
@@ 7.2. Embedding Methodology Comparison
The performance and energy consumption results are illustrated in Figure 10. Both the ME and CE designs demonstrate a dramatic reduction in execution cycles compared to the MA by fully parallelizing the computation. The MA, constrained by the need to fetch weights from SRAM and its limited multiplier array, requires significantly more cycles to complete the same task. The energy reduction of ME is also significant. It consumes the least energy by completely eliminating memory access. While the CE also eliminates power from SRAM access, its massive area leads to substantial leakage power, making it less energy-efficient than ME. The energy consumption of MA is mainly driven by repeated, power-intensive accesses to its SRAM.
@@ 7.3. System-Level Performance Comparison
    WSE-3 data is obtained from published reports (Kundu et al., 2025; Moss, 2025; Wang, 2024; cer, 2024b) and calibrated against performance on its public cloud service (cer, 2024a).
This section provides a comparison of the system-level performance and efficiency of the HNLPU, NVIDIA H100, and Cerebras WSE-3 on gpt-oss 120 B model, as detailed in Table 2. The HNLPU demonstrates orders-of-magnitude advantages in both throughput and energy efficiency, achieving up to $5,555\times$ and $85\times$ throughput, and $1047\times$ and $283\times$ energy efficiency, respectively. The HNLPU’s superior performance and efficiency stem from a fundamental architectural redesign that diverges from conventional systems.
@@
Finally, the HNLPU operates on a highly optimized, model-specific dataflow, workload partitioning, and pipelining strategy, which contrasts with the instruction-driven paradigm of GPUs. This eliminates the significant overhead from control unit such as instruction decoding, scheduling, and control flow. It ensures that nearly all time, power, and area are dedicated to effective computation.
7.4. Economy and Carbon Footprint
Table 3. Total Cost of Ownership (TCO) Analysis for Large-Scale LLM Inference over a 3-Year Lifecycle
Parameter Category Single HNLPU Server 10,000-GPU H100 cluster Remarks / Calculation Basis
@@ 7.4. Economy and Carbon Footprint
[…]
IT Power Load ($\mathrm{MW}$) 0.0552 13 H100: 10k GPUs $\times$ 1.3kW.
Total Datacenter Power ($\mathrm{MW}$) 0.0773 18.2 Power Usage Effectiveness (PUE) = 1.4.
@@
Chips NRE / GPU Price \$ 184 M \$ 300 M Photomasks = \$ 64.6 M, Wafer, Test, Packaging, IP, Tools, Services = \$ 119.4 M
Server / Networking Cost \$ 2 M \$ 150 M Estimated as 50% of GPU price.
Datacenter Infra. Cost \$ 0.04 M \$ 35 M
Total Initial CapEx \$ 186 M \$ 485 M Sum of all CapEx components above.
Annual Update Re-spin Cost \$ 44.3 M 0 Photomasks = \$ 36.9 M, Wafer, Test, Packaging = \$ 7.4 M
@@
Electricity Cost (@ \$0.095/kWh) \$0.19M \$45.44M Total Power × 3 years × \$0.095/kWh (Massed Compute, 2024).
@@
Static Model (No Updates) \$ 186.2 M \$ 530.4 M Total CapEx + 3-Year OpEx.
Dynamic Model (Annual Updates) \$ 274.8 M \$ 530.4 M Static TCO + 2× Update Re-spin.
Performance-Cost Efficiency (Relative)
[…]
@@
Total tCO2e Emissions (Static / Dynamic) 780 / 794 182,321 Following  (Gupta et al., 2022; Zhao et al., 2024b; Li et al., 2025), including embodied and operational emissions.
To evaluate the economic feasibility of the HNLPU in real-world deployments, this section presents a Total Cost of Ownership (TCO) analysis. We compare a single 32 U rack containing eight HNLPU systems against a commercial cluster of 10,000 NVIDIA H100 GPUs delivering equivalent inference throughput.
For Capital Expenditure (CapEx), the HNLPU’s total CapEx is significantly lower than the total investment required for a 10,000-GPU H100 cluster, despite a substantial Non-Recurring Engineering (NRE) cost of \$184 M.
In terms of Operational Expenditure (OpEx), HNLPU’s total electricity cost over three years is only \$0.19 M. This represents an energy cost reduction of over \$45 M compared to the GPU cluster, enabling green and sustainable AI computation.
Overall, within a three-year lifecycle, the HNLPU’s TCO is nearly 65% lower than that of the GPU cluster. Our analysis explicitly accounts for model updating by including the cost of metal-embedding re-spins. In the demanding scenario of yearly model updates, the HNLPU’s TCO maintains a significant advantage.
@@ 8. Discussion
In Section 7.4, we assume the fabrication of only a single HNLPU, as a single unit is already sufficient to meet the demands of today’s high-volume LLM inference services. We anticipate that the unprecedented performance of HNLPU will unlock novel LLM application scenarios that were impossible otherwise, thereby stimulating further growth in inference volume. If multiple systems are produced, NRE will be further amortized, amplifying the cost advantages.
We estimate the initial NRE cost on making HNLPU chips for various LLMs other than gpt-oss in Table 4. Results suggest that a wide range of model sizes can be deployed within an acceptable budget.
Future works extending HNLPU include: (1) Design for Process Variation Tolerance (PVT), adding hardware redundancy and online integrity self-correcting such as Error-Correction Codes (ECC) for higher yield; (2) Enhanced Flexibility on Sea-of-Neurons, enabling hyper-parameter updates with annual re-spin by programmable dataflow; (3) Automated Design and Test, including an automated Hardwired-Neuron Compiler for shorten delay in the design flow; (4) Extended Application Scenarios, implementing LoRA (leaving a small portion of dynamic weights), conditional decoding (programmable sampling algorithms), and support of use cases other than generation (sequence scoring, text-embedding, etc.). (5) Scalability for Extremely Large Models, However, we see no obvious technical obstacles towards implementing these features on HNLPU.
@@
Kimi-K2 (Team et al., 2025) DeepSeek-V3 (Liu et al., 2024) QwQ (Qwen, 2025) Llama-3 (Dubey et al., 2024)
@@ 9. Related Works
This work is mostly inspired by the current progresses in large language models. The success of GPT-3 has proven that LLMs are few-shot learners (Brown et al., 2020). DeepSeek-V3 (Zhao et al., 2025) proved the great benefits from specialization focusing on a single LLM. However, as a software company they have to build upon existing hardware in the market. The mortal computing argument from Hinton (Hinton, 2023) also partly parallels our vision on extremely specialized LPUs.
The first principle of DNN acceleration has been time-multiplexing hardware neurons and capturing data reuse, a wide belief since DianNao in 2014 (Chen et al., 2014). However, the data reuse chances are evaporating from modern LLM inference, which only has $\sim$1 operational intensity in the autoregressive decoding process. Most researchers agree that external memory accesses are the key challenge, while having divergent vision of future. Previous efforts focus on scheduling (Yu et al., 2022; Cai et al., 2025; Mei et al., 2025; Cheng et al., 2025; Sankaralingam et al., 2022; Stojkovic et al., 2025), quantization (Darvish Rouhani et al., 2023; Liu et al., 2025a; Qin et al., 2025; Gil et al., 2025; Lascorz et al., 2024), sparsity (Gondimalla et al., 2023; Wang et al., 2024, 2022; Zhao et al., 2024a), speculative decoding (Kim et al., 2024), and revisiting spiking neurons (Kim et al., 2024), but these techniques cannot fundamentally solve the memory access issue. The Process-In-Memory (PIM) architecture has attracted considerable attention owing to its ability to reduce data movement and alleviate external memory accesses (He et al., 2025; Yue et al., 2025; Heo et al., 2024; Gu et al., 2025; Liu et al., 2025b). However, it cannot fundamentally eliminate the memory accesses required for weight loading, or suffer from high AD/DA overhead.
After the success of LLM, Language Processing Units (LPU) are emerging as the next important processor scheme. Groq released the first commercialized LPU, rebranded from its Tensor Streaming Multiprocessor in 2022 (Abts et al., 2022), which attempts to address the issue by incorporating huge on-chip SRAM. Etched Sohu (etc, 2025) is the most aggressively specialized LPU project that etched Transformer model architecture directly in the fabric. However, these LPUs failed to etch weights, possibly due to the density and cost limitation explained in Section 2. Proposed by this paper, we expect that Metal-Embedding will be the prerequisite technology for the emerging fully-specialized LPU products.
Metal-Embedding shares some underlying techniques with existing architectures: Primitivization architecture proposed by Cambricon-C (Chen et al., 2024), bit-serial architecture explored by Stripes (Judd et al., 2016), and the prefabricated wafers once emerged as sea-of-gates architecture in 1990s (de Lima and Kinniment, 1995).

### v2

@@ 6.1. Hardware and System-Level Evaluation
We implemented the core components of HNLPU architecture, including HN Array, Control Unit, VEX, Interconnection Engine and on-chip Attention Buffer in RTL with Verilog, and verified the correctness of the RTL design using extensive test cases. We followed a standard ASIC design flow to obtain physical characteristics. The design was synthesized using Synopsys Design Compiler and placed-and-routed using Synopsys IC Compiler, on 5 nm technology. Power consumption was analyzed by PrimeTime PX using workload-derived switching activity (SAIF file) to accurately model both static and dynamic power. On-chip SRAMs were generated and analyzed using Memory Compiler on the same technology node.
@@ 6.2. Model
We selected the OpenAI gpt-oss 120 B model for system-level evaluation. It is a state-of-the-art open-source MoE large language model built on Llama-style architecture. We used 4-bit quantized version of the model and hardwiring the weights in HN Array. HNLPU implementation of the model follows partitioning method, dataflow and mapping strategies detailed in Section 4 and Section 5.
@@ 6.3. Baseline Configurations
    Cerebras WSE-3: The throughput was empirically measured through publicly accessible Cerebras cloud service (Cerebras, 2024a) running the gpt-oss 120 B model. As power measurement on cloud is not practical, we adopted the system power figures reported in (Kundu et al., 2025) instead.
@@
    HNLPU: We utilize post-PnR simulations capturing physical layout parasitics and wire delays. This approach provides high-fidelity performance projection, as HNLPU operates on a deterministic Token-In-Token-Out execution model free from software-stack variability.
@@ 7.1. Layout Characteristics
To validate the physical feasibility of HNLPU, we conducted a sign-off-grade implementation flow across representative PVT corners. The design achieves timing closure at 1.0 GHz under worst-case conditions (SSG, 0.675 V, 125 °C), ensuring robust operation under extreme process variations and voltage drops. The design achieves a congestion-free layout with zero overflow. The routing density on ME layers (M8–M11) remains below 70% (lower than typical accelerators), validating the feasibility of ME strategy. Signal integrity is confirmed by parasitic extraction (avg. $R=164\,\Omega$, $C=7.8$ fF), showing manageable coupling effects. Thermal analysis confirms that the power density (avg. 0.3 W/mm2, peak 1.4 W/mm2) is well within the cooling limits of 2.5D packaging. Finally, the layout is DRC/LVS clean, and yield estimation based on Murphy’s model (defect rate 0.11/cm2) confirms the manufacturability of the design.
Table  1 presents the area and power breakdown of a single chip in HNLPU. The HN Array and the Attention Buffer are the dominant components in terms of both area and power. The chip occupies a total area of 827.08 mm2 and has a power consumption of 308.39 W.
The Attention Buffer sustains 80 TB/s bandwidth and 3-cycle latency under worst-case PVT conditions, confirming sufficient performance margins.
@@ 7.2. Embedding Methodology Comparison
The performance and energy consumption results are illustrated in Figure 13. Both the ME and CE designs demonstrate a dramatic reduction in execution cycles compared to the MA by fully parallelizing the computation. Constrained by the need to fetch weights from SRAM and its limited multiplier array, the MA requires significantly more cycles to complete the same task. The energy reduction of ME is also significant. It consumes the least energy by completely eliminating memory access. While the CE also eliminates power from SRAM access, its massive area leads to substantial leakage power, making it less energy-efficient than ME. The energy consumption of MA is mainly driven by repeated, power-intensive accesses to its SRAM.
@@ 7.3. System-Level Performance Comparison
    WSE-3 data is obtained from published reports (Kundu et al., 2025; Moss, 2025; Wang, 2024; Cerebras, 2024b) and calibrated against performance on its public cloud service (Cerebras, 2024a).
This section provides a comparison of the system-level performance and efficiency of HNLPU, NVIDIA H100, and Cerebras WSE-3 on gpt-oss 120 B model, as detailed in Table 2. HNLPU demonstrates orders-of-magnitude advantages in both throughput and energy efficiency, achieving up to $5{,}555\times$ and $85\times$ throughput, and $1{,}047\times$ and $283\times$ energy efficiency, respectively. HNLPU’s superior performance and efficiency stem from a fundamental architectural redesign that diverges from conventional systems.
@@
Finally, HNLPU operates on a highly optimized, model-specific dataflow, workload partitioning, and pipelining strategy, which contrasts with the instruction-driven paradigm of GPUs. This eliminates the significant overhead from control unit such as instruction decoding, scheduling, and control flow. It ensures that nearly all time, power, and area are dedicated to effective computation.
7.4. Execution Time Analysis
Figure 14. Execution time breakdown across varying context lengths. The total execution time is decomposed into inter-chip CXL communication, projection, non-linear operations, attention computation and memory access stalls.
Figure 14 presents the execution time breakdown across varying context lengths. Memory access latency is effectively hidden by the double-buffering mechanism: stalls remain negligible up to 256K tokens, and reach 10.7% at an extreme context length of 512K, where KV cache is loaded from off-chip HBM. In terms of breakdown, the highly optimized computing components expose inter-chip communication as the dominant factor at shorter lengths, while attention computation becomes dominant as the sequence length increases.
7.5. Economic Analysis and Carbon Footprint
Table 3. Total Cost of Ownership (TCO) Analysis for LLM Inference over a 3-Year Lifecycle.
Parameter Category Low Volume High Volume
HNLPU H100 HNLPU H100
@@ 7.5. Economic Analysis and Carbon Footprint
Number of Systems / GPUs1 1 2,000 50 100,000
Total Datacenter Power ($\mathrm{MW}$)2 0.010 3.64 0.483 182
@@
Node Price3 \$ 59.25 M $\sim$ 123.3 M \$ 79.99 M \$ 62.83 M $\sim$ 129.9 M \$ 4,000 M
Data Center Infrastructure4 \$ 0.2100 M \$ 54.93 M \$ 10.30 M \$ 2,747 M
Total Initial CapEx \$ 59.46 M $\sim$ 123.5 M \$ 134.9 M \$ 73.13 M $\sim$ 140.2 M \$ 6,747 M
Update Re-spin Cost5 \$ 18.53 M $\sim$ 37.06 M \$ 0.00 \$ 22.11 M $\sim$ 43.68 M \$ 0.00
@@
Electricity Cost6 \$ 0.0250 M \$ 9.088 M \$ 1.206 M \$ 454.4 M
Maintenance & Support7 \$ 0.0730 M $\sim$ 0.1353 M \$ 47.24 M \$ 0.3650 M $\sim$ 0.6765 M \$ 2,362 M
@@
Static Model (No Updates) \$ 59.56 M $\sim$ 123.7 M \$ 191.2 M \$ 74.70 M $\sim$ 142.1 M \$ 9,563 M
Dynamic Model (Annual Updates) \$ 96.62 M $\sim$ 197.8 M \$ 191.2 M \$ 118.9 M $\sim$ 229.4 M \$ 9,563 M
@@
Total Emissions (tCO2e) (Static / Dynamic)8 102.0 / 106.0 36,600 4,924 / 5,124 1,830,000
  • All figures are rounded to four significant figures. Appendix B presents the detailed assumptions and source references.
We present a comprehensive Total Cost of Ownership (TCO) analysis over a three-year lifecycle in Table 3. We compare HNLPU against an equivalently provisioned NVIDIA H100 GPU cluster delivering comparable inference throughput. We consider two representative deployment volumes: a low-volume deployment corresponding to a single HNLPU node, and a high-volume scenario corresponding to an OpenAI-scale deployment (OpenAI, 2024b, a). We provide both optimistic and pessimistic estimates to account for the sensitivity of key assumptions. For detailed assumptions and source references, please refer to Appendix B.
@@ 8. Discussion
Inference Volume. Section 7.5 analyzed low (single node volume) and high (50 nodes, OpenAI-scale) volumes. These volume settings are based on existing businesses. We anticipate that the unprecedented performance of HNLPU will unlock novel LLM application scenarios that were previously infeasible, thereby stimulating further growth in inference volume. As production volume increases, NRE costs are further amortized, amplifying the cost advantages.
Field-programmable vs Metal-programmable. 1) As the Sea-of-Neurons architecture reduces the weight update re-spin cost to a minor fraction of the TCO, we expect no strong interest towards field-programmable architecture. 2) Introducing area overhead (more chips) to implement dynamic routing would put even more pressure on the dominant bottleneck of the multi-chip interconnection (Figure 14). Advanced interconnection technology (e.g., wafer-scale integration) would put both HNLPU and field-programmable LPU in a stronger position.
Scalability. We estimate the initial NRE cost on making HNLPU chips for various LLMs other than gpt-oss in Table 4. Results suggest that a wide range of model sizes can be deployed within an acceptable budget.
@@
Kimi-K2 (Team et al., 2025) DeepSeek-V3 (Liu et al., 2024) QwQ (Team, 2025) Llama-3 (Dubey et al., 2024)
@@
[…]
Model Updates. HNLPU updates are performance steps. There is no task that GPT-5.2 can handle but GPT-5.1 cannot attempt, just as the release of B100 did not render H100 obsolete. The ”blue-green” deployment model can be adopted for seamless updates: When a model update is validated on GPU testbeds, new ”green” HNLPU can be manufactured while the ”blue” HNLPU continue serving traffic. Estimated turnaround time is 6–8 weeks. The cost is comparable with regular processor re-spins thanks to the Sea-of-Neurons.
Future works. (1) Enhanced Flexibility on Sea-of-Neurons, enabling hyper-parameter updates with annual re-spin by programmable dataflow; (2) Automated Design and Test, including an automated Hardwired-Neuron Compiler for shortening the delay in the design flow; (3) Extended Application Scenarios, implementing conditional decoding (programmable sampling algorithms), and support of use cases other than generation (sequence scoring, text-embedding, etc.); (4) LoRA for Post-Deployment Updates, adding $\sim$1% field-programmable HNs at side-channel to accommodate dynamic weights. However, we foresee no significant technical obstacles to implementing these features on the HNLPU.
@@ 9. Related Works
This work is mostly inspired by the current progress in large language models. The success of GPT-3 has proven that LLMs are few-shot learners (Brown et al., 2020). DeepSeek-V3 (Zhao et al., 2025) proved the great benefits from specialization focusing on a single LLM. However, as a software company they have to build upon existing hardware in the market. The mortal computing argument from Hinton (Hinton, 2023) also partly parallels our vision on extremely specialized LPUs.
The first principle of DNN acceleration has been time-multiplexing hardware neurons and capturing data reuse, a widely held belief since DianNao in 2014 (Chen et al., 2014). However, the data reuse chances are evaporating from modern LLM inference, which only has $\sim$1 operational intensity in the autoregressive decoding process. Most researchers agree that external memory accesses are the key challenge, while having divergent visions of future. Previous efforts focus on scheduling (Yu et al., 2022; Cai et al., 2025; Mei et al., 2025; Cheng et al., 2025; Sankaralingam et al., 2022; Stojkovic et al., 2025), quantization (Darvish Rouhani et al., 2023; Liu et al., 2025b; Qin et al., 2025; Gil et al., 2025; Lascorz et al., 2024), sparsity (Gondimalla et al., 2023; Wang et al., 2024, 2022; Zhao et al., 2024a), speculative decoding (Leviathan et al., 2023), and revisiting spiking neurons (Kim et al., 2024), but these techniques cannot fundamentally solve the memory access issue. The Process-In-Memory (PIM) architecture has attracted considerable attention owing to its ability to reduce data movement and alleviate external memory accesses (He et al., 2025; Yue et al., 2025; Heo et al., 2024; Gu et al., 2025; Liu et al., 2025a). However, it cannot fundamentally eliminate the memory accesses required for weight loading, or suffers from high AD/DA overhead.
After the success of LLM, Language Processing Units (LPU) are emerging as the next important processor scheme. Groq released the first commercialized LPU, rebranded from its Tensor Streaming Multiprocessor in 2022 (Abts et al., 2022), which attempts to address the issue by incorporating huge on-chip SRAM. Etched Sohu (Wikipedia contributors, 2025) is the most aggressively specialized LPU project that etched Transformer model architecture directly in the fabric. However, these LPUs failed to etch weights, possibly due to the density and cost limitations explained in Section 2. Proposed by this paper, we expect that Metal-Embedding will be the prerequisite technology for the emerging fully-specialized LPU products.
Metal-Embedding shares some underlying techniques with existing architectures: Bespoke ML inference fabrics Primitivization architecture proposed by Cambricon-C (Chen et al., 2024), bespoke machine learning accelerators in printed and flexible electronics (Mubarik et al., 2020; Weller et al., 2021; Ozer et al., 2023; Chakraborty, 2024), bit-serial architecture explored by Stripes (Judd et al., 2016), and the prefabricated wafers once emerged as sea-of-gates architecture in 1990s (de Lima and Kinniment, 1995).
@@ 10. Conclusions
Appendix A Detailed Dataflow Description
A.1. Grouped-Query Attention
Figure 10.(II) illustrates the computation dataflow of the 64 query heads projection. The token activation $X$ in the chip array has shape $(1,2880)$. The activation $X$ in all chips is split into four equal slices of $(1,720)$ and each chip takes one of the slices to calculate the partial sum of the query tensor. Each chip contains a private, hard-wired slice of $W_{q}$ with shape $(720,1024)$ inside the HN array. Therefore the product is generated locally and without any weight fetch. The four partial products are summed by a column-internal reduce operation, yielding a 16-head query vector $(1,1024)$. The 16 query heads are then reshaped into $(2,8,64)$, reflecting the Grouped-Query Attention structure where every 8 query heads correspond to a single KV head.
Figure 10.(III) follows a similar spatial pattern for the key path. Each chip multiplies its $(1,720)$ input slice with its hard-wired $W_{k}$ slice $(720,128)$ and emits a partial key vector $(1,128)$. After the column-internal reduction and reshape, the new key head of $(2,64)$ is held in chip# (# $\ell\bmod 4$, where $\ell$ is the token’s position in the sequence). At this point the data layout is intentionally asymmetric: the query vector is fully replicated across columns, whereas each key vector is unique per chip in the sequence dimension, a choice that minimizes the traffic of the subsequent dot-product. The data flow of $X\cdot W_{v}$ mirrors that of $W_{k}$.
Figure 10.(IV) shows the computation of attention weight $Z$ in column-$i$ chips. Every chip already has the complete duplicate of the column-$i$ $Q$ heads $(2,8,64)$. The key tensor is partitioned horizontally, each chip keeps cached $K$ with shape $(2,64,\ell/4)$ in its local KV-cache ($\ell$ is the current context length). The VEX unit multiplies $Q$ head $(2,8,64)$ with $K$ $(2,64,\ell/4)$ and produces the local attention weight $Z$ $(2,8,\ell/4)$. Because each chip sees only $\ell/4$ tokens, a column-wise all-reduce needs to be performed, after which every chip completes the normalization of its local fragment.
Figure 10.(VI) depicts the output projection and first residual path. After the attention score computation, each of the four chips in the same column now holds the same, flattened attention score for 16 heads. The $W_{o}$ matrix is partitioned row-wise across the column group chips, with each column assigned a weight shape of $(1024,2880)$ and each individual chip’s HN array receiving a $(1024,720)$. Each chip generates a partial-$O$ of shape $(1,720)$. These partial-$O$s are combined via one row-wise all-reduce and one column-wise all-gather to yield the final $X_{o}$ of shape (1, 2880) in all chips.
A.2. Feed-Forward Network with MoE
Figure 10.(VII) shows the experts router stage. After the Group Query Attention computation, all chips in HNLPU contain the complete $(1,2880)$ $X_{o}$ vector. These values are passed through RMSNorm before entering the routing layer. As $W_{\text{rout}}$ only accounts for about $0.01\%$ of the total weights, we replicated a copy of the router weights on all 16 chips to avoid inter-chip data exchange. After the computation of $X_{\text{norm}}\cdot W_{\text{rout}}$, each chip obtains complete $X_{\text{rout}}$. Next, top-k and softmax operations are performed. The top-k result is used to generate a masked input tensor, $X$, with a shape of $(128,1,2880)$. In this tensor, the values for the top-k experts are set to those of $X_{\text{norm}}$, while all others are set to zero. Additionally, the top-k results are normalized with a softmax to obtain the expert weights.
Figure 10.(VIII) shows the up- and gate-projection stage. Following the top-k masking, each chip processes its masked $X_{\text{mask}}$ tensor, which consists of 8 vectors, each of shape $(1,2880)$. Among the 128 $X_{\text{mask}}$ $(1,2880)$ vectors in all chips, only $k$ are non-zero. The weight of $W_{\text{up}}$ is sliced into 16 tiles, with the shape of $(8,2880,2880)$ for each chip. Then, the local 8 vectors of $X_{\text{mask}}$ are multiplied with the sliced weight $W_{\text{up}}$ of eight experts. The 16 chips produce a total $X_{\text{up}}$ tensor of shape $(128,1,2880)$. Because each chip stores the complete weight matrices for all experts, this step requires no inter-chip data communication. The gate projection follows the same partitioning: $X_{\text{{mask}}}$ is multiplied by the corresponding $W_{\text{gate}}$ slice, yielding an $X_{G}$, $(128,1,2880)$ in total and $(8,1,2880)$ for each chip. Applying the SwiGLU activation to $X_{G}$ and $X_{\text{{up}}}$, and taking the element-wise product, we get the output $X_{t}$ for the subsequent down-projection.
Figure 10.IX shows the down-projection and second residual path. Still, $X_{t}$ in each chip with the shape of (8, 1, 2880)(8,\,1,\,2880) multiplies its down weight slice $W_{\text{down}}$, (8, 2880, 2880)(8,\,2880,\,2880), to produce a partial $X_{\text{down}}$, (8, 1, 2880)(8,\,1,\,2880). Next, the expert weights, which were obtained from the previous stage (VII), are multiplied with the corresponding expert outputs to get the weighted output for each expert. Subsequently, an all-chip all-reduce operation is performed to sum the partial results. The shape of $X_{\text{down}}$ in all chips is from (128, 1, 2880)(128,\,1,\,2880) to (1, 2880)(1,\,2880). The summed $X_{\text{down}}$ is then added to $X_{o}$ to yield the final layer output Y(1, 2880)Y\,(1,\,2880).
Appendix B Notes to Table 3.
1Deployment scale. We define the ”Low Volume” scenario as a single HNLPU system. The ”High Volume” scenario targets OpenAI-scale throughput ($\sim$100 M tokens/s (OpenAI, 2024b, a)), corresponding to a 50-system HNLPU cluster. To ensure a fair TCO comparison, we normalize hardware counts based on equivalent inference throughput. Under a high-concurrency workload (1K prefill/1K decode, concurrency 50), the average throughput per H100 GPU in a distributed setting is 1.08 K tokens/s (Clarifai, 2025). Consequently, given the HNLPU’s $\sim$2 M tokens/s throughput under the same workload configuration, we equate one HNLPU system to approximately 2,000 H100 GPUs.
2Facility power modeling / PUE. Facility-level PUE is assumed to be 1.4, consistent with modern hyperscale AI datacenters (Davis, 2024).
Table 5. HNLPU Cost Analysis.
- Cost (\$)
Recurring Cost (\$ / chip)
Wafer 629
Package & Test 111 $\sim$ 185
HBM 1,920 $\sim$ 3,840
System Integration 1,900 $\sim$ 3,800
Non-recurring Cost (\$)
Photomask
      Homogeneous Mask 13.85 M $\sim$ 27.69 M
      Metal-Embedding Mask 18.46 M $\sim$ 36.92 M
Design & Development
      Architecture 1.87 M $\sim$ 3.74 M
      Verification 9.97 M $\sim$ 19.93 M
      Physical 4.80 M $\sim$ 14.41 M
      IP 10.23 M $\sim$ 20.46 M
Total Cost Scenarios (\$)
Initial Build (Full NRE + Recurring)
      1-HNLPU 59.25 M $\sim$ 123.3 M
      50-HNLPU 62.83 M $\sim$ 129.9 M
Re-spin (Metal-Embedding Mask + Recurring)
      1-HNLPU 18.53 M $\sim$ 37.06 M
      50-HNLPU 22.11 M $\sim$ 43.68 M
3Node price of H100 and HNLPU.For H100 price, each NVIDIA HGX H100 platform (8 GPUs/node) costs about \$ 320,000, including server, intra-node networking and 3-year hardware warranty (Thinkmate, 2024; Router-Switch.com, 2024). For HNLPU, we break down the node cost in Table 3 into recurring cost and non-recurring engineering (NRE) cost in Table 5.
Regarding recurring cost, we first estimate the silicon cost. Assuming a cost of \$ 16,988 for a 300 mm 5 nm wafer (Center for Security and Emerging Technologies (CSET), 2020; Khan and Mann, 2020), Murphy’s model ($D_{0}=0.11$ def/cm2) predicts a 43% yield ($\sim$27 of 62 dies), resulting in $\approx\mathdollar 629$ per good die. Second, packaging and testing are estimated at \$3,000–\$5,000 per wafer to account for 2.5D integration complexity (Palesko et al., 2013), resulting in an amortized cost of \$111–\$185 per chip. Third, given an HBM cost of \$10–\$20/GB (EEWorld, 2025; Depend Electronics, 2024), the 8-stack configuration (24 GB per stack) amounts to \$1,920–\$3,840 per HNLPU module. Finally, we include system integration costs, covering the chassis, motherboard, cooling, power, and CXL interconnects; these figures align with per-chip costs from established commercial platforms (Patel and Wong, 2023).
Regarding one-time costs, we distinguish between photomask NRE and design & development expenses. First, we model the photomask NRE using a normalized cost model based on lithography complexity for the 5 nm technology node. To account for the disparity in manufacturing costs, we assign a cost weighting factor of $6\times$ to EUV reticles relative to standard 193i DUV reticles (Scheid, 2023). Given a typical 5 nm layer stack comprising 12 EUV and 58 DUV layers (Nenni, 2025; AnySilicon, 2026; Vashishtha and Clark, 2022; Jones, 2020), the total mask set value corresponds to $58+(12\times 6)=130$ normalized DUV units. In HNLPU architecture, the metal-embedding configuration requires 10 DUV reticles (VIA7, M8 Mandrel, M8 Cut, VIA8, M9 Mandrel, M9 Cut, VIA9, M10, VIA10, M11); consequently, this variable portion accounts for $7.7\%$ ($10/130$) of the full mask set, while the remaining $92.3\%$ represents the homogeneous mask cost shared across all variants. Anchoring the absolute 5 nm mask set cost to a range of \$ 15 M (optimistic) to \$ 30 M (pessimistic(Graening et al., 2024; Patel, 2022; Kasprowicz, 2017), we derive a shared homogeneous mask cost of \$ 13.85–\$ 27.69 M. The metal-embedding cost is estimated at \$ 1.15–\$ 2.31 M per variant, amounting to \$ 18.46–\$ 36.92 M in total for 16 chips. Second, for HNLPU design and development costs, we derive our estimates from internal engineering data and design experience.
4Data center infrastructure. We consider two primary capital expenditures: inter-node networking and facility construction. For the H100 cluster baseline, we assume a standard three-tier non-blocking Fat-Tree topology. In terms of hardware composition, the network fabric is built using NVIDIA ConnectX-7 (Colfax International, 2024) network interface cards (NICs) and Quantum-2 (QM9700) InfiniBand switches (FS.com, 2024b), interconnected with corresponding optical transceivers (FS.com, 2024a). Accounting for the NICs, switches, and cabling costs, the estimated network equipment capital expenditure is approximately \$45 K per node. The facility construction cost is modeled as \$12 M per MW of critical IT load (Cushman & Wakefield, 2025). For HNLPU, we scale the networking cost based on the number of chips, while the construction cost is scaled based on total power consumption.
5Re-spin cost. For H100, changing the model does not require a re-spin; therefore, this cost is zero. For HNLPU, the re-spin cost comprises the Non-Recurring cost for the metal-embedding mask, plus the recurring costs for fabricating, packaging, and testing, as listed in Table 5.
6Electricity. The electricity cost is calculated based on the industrial electricity price of \$0.095/kWh, representative of major U.S. data center hubs (U.S. Energy Information Administration, 2024).
7Maintenance & Support. For H100 clusters, Maintenance & Support includes software licensing and hardware maintenance. Software license fees are calculated based on NVIDIA AI Enterprise pricing guidelines (NVIDIA, 2024). Hardware maintenance is conservatively estimated as 5% of hardware CapEx per year (Rasmussen, n.d.; Koomey et al., 2007). For HNLPU, we model this cost by provisioning spare nodes: one for the low-volume scenario and five for the high-volume scenario.
8CO2 emission. Our carbon emission analysis incorporates both embodied emissions from hardware manufacturing and operational emissions from energy consumption. We estimate the manufacturing emission for a single H100 card or an HNLPU module at 124.9 kgCO2(Zhao et al., 2024b; Li et al., 2025). The grid carbon intensity is assumed to be 0.38 kgCO2e per kWh (Gupta et al., 2022).